# Wen Wang    Ruibing Hou    Hong Chang    Shiguang Shan    Xilin Chen

### v2

@@ 5 Conclusion
Limitations. MATS is currently focused on the audio domain. In the future, we intend to extand MATS by incorporating additional modalities, such as visual data and 3D point clouds, aiming to develop a more versatile multimodal model capable of handling a wider range of modalities, all while relying exclusively on text-based supervision.
@@
$\begin{array}[]{ccc}R_{\mathrm{te}}(h)&=&R_{\mathrm{te}}(h)-R_{\mathrm{tr}}(h)+R_{\mathrm{tr}}(h)-\hat{R}_{\mathrm{tr}}(h)+\hat{R}_{\mathrm{tr}}(h)\\ &&\\ &\leq&\big{|}R_{\mathrm{te}}(h)-R_{\mathrm{tr}}(h)\big{|}+\big{|}R_{\mathrm{tr}}(h)-\hat{R}_{\mathrm{tr}}(h)\big{|}+\hat{R}_{\mathrm{tr}}(h)\\ \end{array}$ (8)
@@ B.1 The train dataset of MATS-GPT2 and MATS-LLaMA
RAVDESS (Livingstone & Russo, 2018)
CAP WavCaps (Mei et al., 2024), Macs (Martín-Morató & Mesaros, 2021), $\sim$518K $\sim$34.5%
@@ B.2 The train dataset of MATS-Audio
CAP WavCaps (Mei et al., 2024), Macs (Martín-Morató & Mesaros, 2021), $\sim$517K $\sim$34.2%
@@ B.3 The test benchmark of MATS
[…]
@@
[…]
@@
In this part, we describe the pipeline and prompt used with ChatGPT (Achiam et al., 2023) to generate audio descriptions for various classification task datasets, including ESC-50 (Piczak, 2015), FSD50K (Fonseca et al., 2021), VGG (Chen et al., 2020), ECS-50 (Piczak, 2015), US8K (Salamon et al., 2014), TUT (Mesaros et al., 2016), SESA (Spadini, 2019), VS (Gong et al., 2022), CREMA-D (Cao et al., 2014), and RAVDESS (Livingstone & Russo, 2018) for sound classification, and GTZAN (Sturm, 2014) and BJO (Tian et al., 2014) for music classification. Figure S5 provides an overview of the pipeline. The pipeline involves supplying GPT-4o-mini with a set of example audio descriptions corresponding to a few labels, followed by a prompt to guide the generation process. GPT-4o-mini then produces descriptions for all provided labels. The final set of generated descriptions, paired with their corresponding labels, serves as input for MATS.
@@ C.2 Hyperparameter Analysis
The impact of the different balancing parameter $\lambda$.  In this section, we conducted experiments with different values of $\lambda$, as shown in Figure S6. The results clearly demonstrate that the best performance is achieved when both language and audio are integrated simultaneously, and the performance is relatively insensitive to the variations in $\lambda$. However, when audio is entirely removed, the model’s performance drops significantly, further validating the effectiveness of the balancing strategy within the Santa mechanism.

### v3

@@ 5 Conclusion
Acknowledgements
This work is partially supported by National Natural Science Foundation of China (NSFC): 62376259 and 62306301, National Postdoctoral Program for Innovative Talents under Grant BX20220310.
Impact Statement
This paper presents work whose goal is to advance the field of Large Audio-Language Models. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
[…]
@@ B.1 The train dataset of MATS-GPT2 and MATS-LLaMA
RAVDESS (Livingstone and Russo, 2018)
CAP WavCaps (Mei et al., 2024), Macs (Martín-Morató and Mesaros, 2021), $\sim$518K $\sim$34.5%
@@ B.2 The train dataset of MATS-Audio
CAP WavCaps (Mei et al., 2024), Macs (Martín-Morató and Mesaros, 2021), $\sim$517K $\sim$34.2%
@@ B.3 The test benchmark of MATS
CAP AudioCaps (Kim et al., 2019) $\text{CIDEr}|\text{SPICE}|\text{SPIDEr}$
CAP Clotho-v2 (Drossos et al., 2020) $\text{CIDEr}|\text{SPICE}|\text{SPIDEr}$
AQA Clotho-AQA (Lipping et al., 2022) $\text{ACC}|\text{B-ACC}$
@@
CAP MusicCaps (Agostinelli et al., 2023) $\text{BLUE4}|\text{ROUGH-L}$
@@
In this part, we describe the pipeline and prompt used with ChatGPT (Achiam et al., 2023) to generate audio descriptions for various classification task datasets, including ESC-50 (Piczak, 2015), FSD50K (Fonseca et al., 2021), VGG (Chen et al., 2020), ECS-50 (Piczak, 2015), US8K (Salamon et al., 2014), TUT (Mesaros et al., 2016), SESA (Spadini, 2019), VS (Gong et al., 2022), CREMA-D (Cao et al., 2014), and RAVDESS (Livingstone and Russo, 2018) for sound classification, and GTZAN (Sturm, 2014) and BJO (Tian et al., 2014) for music classification. Figure S4 provides an overview of the pipeline. The pipeline involves supplying GPT-4o-mini with a set of example audio descriptions corresponding to a few labels, followed by a prompt to guide the generation process. GPT-4o-mini then produces descriptions for all provided labels. The final set of generated descriptions, paired with their corresponding labels, serves as input for MATS.
@@ C.2 Hyperparameter Analysis
The impact of the different balancing parameter $\lambda$. In this part, we conducted experiments with different values of $\lambda$, as shown in Figure S5. The results clearly demonstrate that the best performance is achieved when both language and audio are integrated simultaneously, and the performance is relatively insensitive to the variations in $\lambda$. However, when audio is entirely removed, the model’s performance drops significantly, further validating the effectiveness of the balancing strategy within the Santa mechanism.